OpenStudy (anonymous):

check wheather the ordered pair is a soultion of the system (2,-4) 2x+y=0 -x+2y=10

OpenStudy (aroub):

not a solution

OpenStudy (anonymous):

no -(2)+2(-4)=10 -2-8=10 -10=10 not true

OpenStudy (anonymous):

thanks